OpenStudy (anonymous):

use sqrt property to solve: (4k-1)^2+28=1 k=??

jimthompson5910 (jim_thompson5910):

\[\large (4k-1)^2+28=1 \] \[\large (4k-1)^2=1-28 \] \[\large (4k-1)^2=-27 \] \[\large 4k-1=\pm\sqrt{-27}\] \[\large 4k-1=\sqrt{-27} \ \ \textrm{or} \ \ 4k-1=-\sqrt{-27}\] \[\large 4k-1=3i\sqrt{3} \ \ \textrm{or} \ \ 4k-1=-3i\sqrt{3}\] \[\large 4k=1+3i\sqrt{3} \ \ \textrm{or} \ \ 4k=1-3i\sqrt{3}\] \[\large k=\frac{1+3i\sqrt{3}}{4} \ \ \textrm{or} \ \ k=\frac{1-3i\sqrt{3}}{4}\]

OpenStudy (anonymous):

oh hey!! the page didn't show this for me! You rock :)

jimthompson5910 (jim_thompson5910):

yeah the messaging system is a bit weird for me too